# Hopfield-Fenchel-Young Networks: A Unified Framework for Associative Memory Retrieval

### v2

@@ 1 Introduction
Hopfield networks are biologically plausible neural networks with associative memory capabilities (Amari, 1972; Nakano, 1972; Hopfield, 1982). Their attractor dynamics, which describe how the networks evolve toward stable states or patterns, make them suitable for modeling associative memory retrieval in humans and animals (Tyulmankov et al., 2021; Whittington et al., 2021). The limited storage capacity of classical Hopfield networks was recently overcome through the proposal of new energy functions. These energies were initially proposed for dense associative models by Krotov and Hopfield (2016) and Demircigil et al. (2017) and later expanded to continuous states by Ramsauer et al. (2021), resulting in exponential storage capacity and renewed interest in ”modern” Hopfield networks. In particular, Ramsauer et al. (2021) revealed connections to attention layers in transformers via an update rule linked to the convex-concave procedure (CCCP; Yuille and Rangarajan 2003). However, this model only approximates stored patterns, requiring careful temperature tuning to avoid converging to large metastable states that mix multiple input patterns instead of matching a single pattern.
A common element in many recent studies on Hopfield networks is connecting an energy function with a desired update rule, typically by modeling the network’s temporal dynamics through differential equations, which are discretized to produce updates involving derivatives of Lagrangian terms in the energy function (Krotov and Hopfield, 2021). However, a comprehensive framework for formulating the equations that govern the dynamics of the entire spectrum of Hopfield networks is lacking—a gap we aim to fill by explicitly using convex analysis and Fenchel-Young duality as building blocks (Rockafellar, 1970; Bauschke and Combettes, 2017). This not only allows for designing new energy functions but also provides a way to understand functionalities in transformer architectures, like multi-head attention (Vaswani et al., 2017) and layer normalization (Ba et al., 2016), as well as other normalization techniques (Nguyen and Salazar, 2019). Developing a generalized framework is crucial for a unified theoretical basis to understand and extend Hopfield networks. This unified view facilitates comparative analysis, promotes the discovery of underlying principles that govern Hopfield network dynamics, and potentially leads to improvements in associative memory design and functionality.
@@
[…]
@@
§1 Introduction;
§2 Background;
§3 Hopfield-Fenchel-Young Energies;
§4 Sparse Hopfield Networks;
§5 Structured Hopfield Networks;
§6 Mechanics of Memory Retrieval Modeling;
§7 Experiments;
§8 Related Work;
§9 Conclusions.
@@ 2.1 Associative Memories and Hopfield Networks
In associative memories, data patterns are retrieved based on similarity to a given query, rather than through an explicit address. When a noisy cue of the memories is provided as the query, the aim is to retrieve the most similar memory pattern. Hopfield networks (Amari, 1972; Nakano, 1972; Hopfield, 1982) are neural models inspired by statistical physics, specifically by the Ising model (Ising, 1925), which describes a system of magnetic moments or “spins” of particles that can be in one of two states (±$(\pm 1)$, interacting with each other to minimize the system’s overall energy. In a classical Hopfield network (Hopfield, 1982), states are binary and interact through synaptic weights analogous to the spin-spin couplings in the Ising model. When a query is given as input, the network iteratively adjusts to minimize its energy, leading to the emergence of stable states or attractors, which correspond to stored memory patterns. Pioneering works on classic Hopfield networks, as well as subsequent research by Amit et al. (1985a) and Hertz et al. (1991), has demonstrated the ability of this approach to perform tasks like pattern retrieval and completion for binary data.
@@
where $\bm{W}=\bm{X}^{\top}\bm{X}\in\mathbb{R}^{D\times D}$ is a weight matrix and $\bm{q}\in\{\pm 1\}^{D}$ is a binary vector, leading to the update rule $\bm{q}^{(t+1)}=\mathrm{sign}(\bm{W}\bm{q}^{(t)})$. In this classical construction, the stored patterns, as well as the queries, are assumed to be binary vectors in {±$\{\pm 1\}^{D}$. The discreteness of the update rule (due to the sign transformation) makes this network capable of exact retrieval (i.e., it is able to retrieve memorized patterns perfectly, upon convergence). However, its main limitation is that it has only $N=\mathcal{O}(D)$ memory storage capacity. When this capacity is exceeded, the patterns start to interfere (Amit et al., 1985b; McEliece et al., 1987), resulting in the retrieval of metastable states or spurious attractor points.
@@ 2.2 Modern Hopfield Networks
Ramsauer et al. (2021) revealed an interesting relation between the updates in this modern Hopfield network and the attention layers in transformers. Namely, the minimization of the energy (2) using the concave-convex procedure (CCCP; Yuille and Rangarajan 2003) leads to the update rule
@@
When $\beta=\frac{1}{\sqrt{D}}$, each update matches the computation performed in the attention layer of a transformer with a single attention head and identity projection matrices. This triggered interest in developing variants of Hopfield layers which can be used as drop-in replacements for multi-head attention layers (Hoover et al., 2023).
@@ 2.3 Regularized Argmax and Fenchel-Young Losses
Given a strictly convex function $\Omega:\mathbb{R}^{N}\rightarrow\bar{\mathbb{R}}$ with domain $\mathrm{dom}(\Omega)$, the $\Omega$-regularized argmax transformation (Niculae and Blondel, 2017), $\hat{\bm{y}}_{\Omega}:\mathbb{R}^{N}\rightarrow\mathrm{dom}(\Omega)$ is:
$\hat{\bm{y}}_{\Omega}(\bm{\theta}):=\nabla\Omega^{*}(\bm{\theta})=\operatorname*{argmax}_{\bm{y}\in\mathrm{dom}(\Omega)}\bm{\theta}^{\top}\bm{y}-\Omega(\bm{y}).$ (4)
A trivial example of such a regularized argmax function (4) is obtained when $\Omega(\bm{y})=\frac{1}{2}\|\bm{y}\|^{2}$ with $\mathrm{dom}(\Omega)=\mathbb{R}^{N}$, which leads to the identity map $\bm{y}_{\Omega}(\bm{\theta})=\bm{\theta}$. In general, we are interested in cases where $\mathrm{dom}(\Omega)\subsetneq\mathbb{R}^{N}$, for example the probability simplex $\mathrm{dom}(\Omega)=\triangle_{N}$ (studied in §4) or a polytope (studied in §5). A famous instance of the former is the softmax transformation, obtained when the regularizer is the Shannon negentropy, $\Omega(\bm{y})=\sum_{i=1}^{N}y_{i}\log y_{i}+I_{\triangle_{N}}(\bm{y})$. Another instance is the sparsemax transformation, obtained when $\Omega(\bm{y})=\frac{1}{2}\|\bm{y}\|^{2}+I_{\triangle_{N}}(\bm{y})$ (Martins and Astudillo, 2016), and which corresponds to the Euclidean projection onto the probability simplex. In §4, we analyze a wider set of transformations induced by generalized entropies which include these as particular cases.
@@ 3.1 Definition
  • Minimizing $E_{\mathrm{convex}}$ serves as a regularization, encouraging $\bm{q}$ to stay close to $\bm{X}^{\top}\bm{u}$.
@@ 3.3 Particular cases
Classic HNs $\frac{1}{2}\|\bm{y}\|^{2}$ $\mathbb{R}^{N}$ $-\frac{1}{\beta}H_{\mathrm{b}}\left(\frac{\mathbf{1}+\bm{q}}{2}\right)$ $[-1,1]^{D}$ $\tanh(\beta\bm{X}^{\top}\bm{X}\bm{q}^{(t)})$
@@
Poly-DAMs $s^{-1}\|\bm{y}\|_{s}^{s}$ $\mathbb{R}^{N}$ $-\frac{1}{\beta}H_{\mathrm{b}}\left(\frac{\mathbf{1}+\bm{q}}{2}\right)$ $[-1,1]^{D}$ $\tanh(\beta\bm{X}^{\top}\mathrm{spow}(\bm{X}\bm{q}^{(t)},r-1))$
@@
Exp-DAMs $\bm{y}^{\top}(\log\bm{y}-\mathbf{1})_{+}$ $\mathbb{R}_{+}^{N}$ $-\frac{1}{\beta}H_{\mathrm{b}}\left(\frac{\mathbf{1}+\bm{q}}{2}\right)$ $[-1,1]^{D}$ $\tanh(\beta\bm{X}^{\top}\exp(\bm{X}\bm{q}^{(t)})$
@@ 3.3.2 Dense associative memories (DAMs)
This corresponds to choosing $\Omega(\bm{y})=-s^{-1}\|\bm{y}\|_{s}^{s}$, where $r^{-1}+s^{-1}=1$ (.$\|.\|_{r}$ and .$\|.\|_{s}$ are dual norms), and $\Psi(\bm{q})$ as in classical Hopfield networks. We have $(\nabla\Omega^{*})(\bm{\theta})=\mathrm{sign}(\bm{\theta})|\bm{\theta}|^{r-1}=:\mathrm{spow}(\bm{\theta},r-1)$, where spow denotes the signed power function. The update rule is
@@ 4.5 Properties: Margins, sparsity, and exact retrieval
Assume patterns are randomly placed on the sphere with uniform distribution. Then, with probability $1-p$, the HFY network can store and exactly retrieve $N=\mathcal{O}(\sqrt{p}\zeta^{\frac{D-1}{2}})$ patterns in one iteration under a $\epsilon$-perturbation if
@@ 4.6 Extension of previous results for extra normalization step
[…]
@@ 5.4 Guarantees for retrieval of pattern associations
For the sequential $k$-subsets problem in Example 2, we have $r=2N-1$. Noting that any two distinct $\bm{y}$ and $\bm{y}^{\prime}$ differ in at most $2k$ variable nodes, and since each variable node can affect 6 bits (2 for $\bm{y}_{V}$ and 4 for $\bm{y}_{F}$), the Hamming distance between $\bm{y}$ and $\bm{y}^{\prime}$ is at most $12k$, therefore we have $D_{i}=\sqrt{12k}$, which leads to the condition $\Delta_{i}\geq\frac{6k}{\beta}+12Mk\epsilon$.
@@ 6.1 Free recall
Algorithm 1 Free recall with constrained sparsemax. $\bm{X}\in\mathbb{R}^{N\times D}$ represents the memory and $\bm{q}\in\mathbb{R}^{D}$ denotes the query, initialized as an arbitrary cue. $T$ is the number of inner associative Hopfield iterations. $N$ is the number of memory patterns, which equals the number of outer iterations during the recall process.
1:$\bm{X}$, $\bm{q}$, $\beta$, $T$
2:$\bm{u}\leftarrow\mathbf{1}_{N}$triangleright Initialize upper bounds
@@
5:         $\bm{\theta}\leftarrow\bm{X}\bm{q}$triangleright Scores
@@
7:         $\bm{q}\leftarrow\bm{X}^{\top}\bm{p}$triangleright Hopfield update
8:    $\bm{u}\leftarrow\bm{u}-\bm{p}$triangleright Upper bound probabilities
@@
1:$\bm{X}$, $\bm{q}$, $\lambda$, $\tau$, $\beta$, $T$
2:$\bm{a}\leftarrow\mathbf{0}_{N}$triangleright Initial average
@@
4:    $\bm{p}\leftarrow\alpha\text{-entmax}\left(\beta\left(\bm{X}\bm{q}-\lambda\bm{a}\right)\right)$triangleright Penalized probabilities
5:    $\bm{a}\leftarrow\tau\bm{p}+(1-\tau)\bm{a}$triangleright Exponentially weighted average
6:    $\bm{q}\leftarrow\bm{X}^{\top}\bm{p}$triangleright Outer Hopfield update
@@
9:         $\bm{q}\leftarrow\bm{X}^{\top}\bm{p}$triangleright Inner Hopfield update

### v5

@@ 1 Introduction
Hopfield networks are biologically plausible neural networks with associative memory capabilities (Amari, 1972; Nakano, 1972; Hopfield, 1982). Their attractor dynamics, which describe how the networks evolve toward stable states or patterns, make them suitable for modeling associative memory retrieval in humans and animals (Tyulmankov et al., 2021; Whittington et al., 2021). The limited storage capacity of classical Hopfield networks was recently overcome through the proposal of new energy functions. These energies were initially proposed for dense associative models by Krotov and Hopfield (2016) and Demircigil et al. (2017) and later expanded to continuous states by Ramsauer et al. (2021), resulting in exponential storage capacity and renewed interest in ”modern” Hopfield networks. In particular, Ramsauer et al. (2021) revealed connections to attention layers in transformers via an update rule linked to the convex-concave procedure (CCCP; Yuille and Rangarajan 2003). However, this model only approximates stored patterns, requiring careful temperature tuning to avoid converging to large metastable states that mix multiple input patterns.
A common element in many recent studies on Hopfield networks is connecting an energy function with a desired update rule, typically by modeling the network’s temporal dynamics through differential equations, which are discretized to produce updates involving derivatives of Lagrangian terms in the energy function (Krotov and Hopfield, 2021). However, a comprehensive framework for formulating the equations that govern the dynamics of the entire spectrum of Hopfield networks is lacking—a gap we aim to fill by explicitly using convex analysis and Fenchel-Young duality as building blocks (Rockafellar, 1970; Bauschke and Combettes, 2017). This not only allows for designing new energy functions but also provides a way to understand functionalities in transformer architectures, like multi-head attention (Vaswani et al., 2017) and layer normalization (Ba et al., 2016), as well as other normalization techniques (Nguyen and Salazar, 2019). Developing a generalized framework is crucial for a unified theoretical basis to understand and extend Hopfield networks. This unified view facilitates comparative analysis, reveals underlying principles that govern their dynamics, and potentially leads to improvements in associative memory design and functionality.
@@
We denote by $\triangle_{N}$ the $(N-1)\textsuperscript{th}$-dimensional probability simplex, $\triangle_{N}:=\{\bm{p}\in\mathbb{R}^{N}:\bm{p}\geq\mathbf{0},\,\mathbf{1}^{\top}\bm{p}=1\}$. The convex hull of a set $\mathcal{Y}\subseteq\mathbb{R}^{M}$ is $\mathrm{conv}(\mathcal{Y}):=\{\sum_{i=1}^{N}p_{i}\bm{y}_{i}\,:\,\bm{p}\in\triangle_{N},\,\bm{y}_{1},...,\bm{y}_{N}\in\mathcal{Y},\,N\in\mathbb{N}\}$. We have $\triangle_{N}=\mathrm{conv}(\{\bm{e}_{1},...,\bm{e}_{N}\})$, where $\bm{e}_{i}\in\mathbb{R}^{N}$ is the $i$th basis (one-hot) vector. Given a convex function $\Omega:\mathbb{R}^{N}\rightarrow\bar{\mathbb{R}}$, where $\bar{\mathbb{R}}={\mathbb{R}}\cup\{+\infty\}$, we denote its domain by $\mathrm{dom}(\Omega):=\{\bm{y}\in\mathbb{R}^{N}\,:\,\Omega(\bm{y})<+\infty\}$ and its Fenchel conjugate by $\Omega^{*}(\bm{\theta})=\sup_{\bm{y}\in\mathbb{R}^{N}}\bm{y}^{\top}\bm{\theta}-\Omega(\bm{y})$. We denote by $I_{\mathcal{C}}$ the indicator function of a convex set $\mathcal{C}$, defined as $I_{\mathcal{C}}(\bm{y})=0$ if $\bm{y}\in\mathcal{C}$, and $I_{\mathcal{C}}(\bm{y})=+\infty$ otherwise.
@@
§1
Introduction;
§2
Background;
§3
Hopfield-Fenchel-Young Energies;
§4
Sparse Hopfield Networks;
§5
Structured Hopfield Networks;
§6
Mechanics of Memory Retrieval Modeling;
§7
Experiments;
§8
Related Work;
§9
Conclusions.
@@ 2.1 Associative Memories and Hopfield Networks
In associative memories, data patterns are retrieved based on similarity to a given query, rather than through an explicit address. When a noisy cue of the memories is provided as the query, the aim is to retrieve the most similar memory pattern. Hopfield networks (Amari, 1972; Nakano, 1972; Hopfield, 1982) are neural models inspired by statistical physics, specifically by the Ising model (Ising, 1925), which describes a system of magnetic moments or “spins” of particles that can be in one of two states $(\pm 1)$, interacting with each other to minimize the system’s overall energy. In a classical Hopfield network (Hopfield, 1982), states are binary and interact through synaptic weights analogous to the spin-spin couplings in the Ising model. When a query is given as input, the network iteratively adjusts to minimize its energy, leading to the emergence of stable states or attractors, which correspond to stored memory patterns. Pioneering works on classic Hopfield networks, as well as subsequent research by Amit et al. (1985a) and Hertz et al. (1991), has demonstrated the ability of this approach to perform tasks like pattern retrieval and completion for binary data.
@@
where $\bm{W}=\bm{X}^{\top}\bm{X}\in\mathbb{R}^{D\times D}$ is a weight matrix and $\bm{q}\in\{\pm 1\}^{D}$ is a binary vector, leading to the update rule $\bm{q}^{(t+1)}=\mathrm{sign}(\bm{W}\bm{q}^{(t)})$. In this classical construction, the stored patterns, as well as the queries, are assumed to be binary vectors in $\{\pm 1\}^{D}$. The discreteness of the update rule (due to the sign transformation) makes this network capable of exact retrieval (i.e., it is able to retrieve memorized patterns perfectly, upon convergence). However, its main limitation is that it has only $N=\mathcal{O}(D)$ memory storage capacity. When this capacity is exceeded, the patterns start to interfere (Amit et al., 1985b; McEliece et al., 1987), resulting in the retrieval of metastable states or spurious attractor points.
@@ 2.2 Modern Hopfield Networks
Ramsauer et al. (2021) revealed an interesting relation between the updates in this modern Hopfield network and the attention layers in transformers. Namely, the minimization of the energy (2) using the concave-convex procedure (CCCP; Yuille and Rangarajan 2003), also known as difference-of-convex (DC) programming (Horst and Thoai, 1999), leads to the update rule
@@
When $\beta=\frac{1}{\sqrt{D}}$, each update matches the computation performed in the attention layer of a transformer with a single attention head and identity projection matrices (Ramsauer et al., 2021, §2). This triggered interest in developing variants of Hopfield layers which can be used as drop-in replacements for multi-head attention layers (Hoover et al., 2023).
@@ 2.3 Regularized Argmax and Fenchel-Young Losses
Given a strictly convex function $\Omega:\mathbb{R}^{N}\rightarrow\bar{\mathbb{R}}$ with compact domain $\mathrm{dom}(\Omega)$, the $\Omega$-regularized argmax transformation (Niculae and Blondel, 2017), $\hat{\bm{y}}_{\Omega}:\mathbb{R}^{N}\rightarrow\mathrm{dom}(\Omega)$ is:
$\hat{\bm{y}}_{\Omega}(\bm{\theta}):=\nabla\Omega^{*}(\bm{\theta})=\operatorname*{argmax}_{\bm{y}\in\mathrm{dom}(\Omega)}\bm{\theta}^{\top}\bm{y}-\Omega(\bm{y}),$ (4)
where $\nabla\Omega^{*}(\bm{\theta})$ denotes the gradient of the Fenchel conjugate $\Omega^{*}$ at $\bm{\theta}$. [^2] Note that $\Omega^{*}$ is guaranteed to be everywhere differentiable under the assumption that $\mathrm{dom}(\Omega)$ is compact and $\Omega$ is strictly convex, which together ensure existence and uniqueness of a maximum in (4). A trivial example of such a regularized argmax function (4) is obtained when $\Omega(\bm{y})=\frac{1}{2}\|\bm{y}\|^{2}$ with $\mathrm{dom}(\Omega)=\mathbb{R}^{N}$, which leads to the identity map $\bm{y}_{\Omega}(\bm{\theta})=\bm{\theta}$. In general, we are interested in cases where $\mathrm{dom}(\Omega)\subsetneq\mathbb{R}^{N}$, for example the probability simplex $\mathrm{dom}(\Omega)=\triangle_{N}$ (studied in §4) or a polytope (studied in §5). A famous instance of the former is the softmax transformation, obtained when the regularizer is the Shannon negentropy, $\Omega(\bm{y})=\sum_{i=1}^{N}y_{i}\log y_{i}+I_{\triangle_{N}}(\bm{y})$. Another instance is the sparsemax transformation, obtained when $\Omega(\bm{y})=\frac{1}{2}\|\bm{y}\|^{2}+I_{\triangle_{N}}(\bm{y})$ (Martins and Astudillo, 2016), and which corresponds to the Euclidean projection onto the probability simplex. In §4, we analyze a wider set of transformations induced by generalized entropies which include these as particular cases.
@@ 3.1 Definition
  • Minimizing $E_{\mathrm{convex}}$ serves as a regularization, encouraging $\bm{q}$ to stay close to $\bm{X}^{\top}\bm{u}$. More intuition about this regularization will be provided in §4.4.
@@ 3.3 Particular cases
Classic HNs $\frac{1}{2}\|\bm{y}\|^{2}$ $\mathbb{R}^{N}$ $\frac{1}{\beta}H_{\mathrm{b}}\left(\frac{\mathbf{1}+\bm{q}}{2}\right)$ $1,1]^{D}$ $\tanh(\beta\bm{X}^{\top}\bm{X}\bm{q}^{(t)})$
@@
Poly-DAMs $s^{-1}\|\bm{y}\|_{s}^{s}$ $\mathbb{R}^{N}$ $\frac{1}{\beta}H_{\mathrm{b}}\left(\frac{\mathbf{1}+\bm{q}}{2}\right)$ $1,1]^{D}$ $\tanh(\beta\bm{X}^{\top}\mathrm{spow}(\bm{X}\bm{q}^{(t)},r-1))$
@@
Exp-DAMs $\bm{y}^{\top}(\log\bm{y}-\mathbf{1})_{+}$ $\mathbb{R}_{+}^{N}$ $\frac{1}{\beta}H_{\mathrm{b}}\left(\frac{\mathbf{1}+\bm{q}}{2}\right)$ $1,1]^{D}$ $\tanh(\beta\bm{X}^{\top}\exp(\bm{X}\bm{q}^{(t)})$
@@ 3.3.2 Dense associative memories (DAMs)
This corresponds to choosing $\Omega(\bm{y})=-s^{-1}\|\bm{y}\|_{s}^{s}$, where $r^{-1}+s^{-1}=1$ ($\|.\|_{r}$ and $\|.\|_{s}$ are dual norms), and $\Psi(\bm{q})$ as in classical Hopfield networks. We have $(\nabla\Omega^{*})(\bm{\theta})=\mathrm{sign}(\bm{\theta})|\bm{\theta}|^{r-1}=:\mathrm{spow}(\bm{\theta},r-1)$, where spow denotes the signed power function. The update rule is
@@ 4.5 Properties: Margins, sparsity, and exact retrieval
Assume patterns are randomly placed on the sphere with uniform distribution. Then, with probability 1p1-p, the HFY network can store and exactly retrieve $N=\mathcal{O}(\sqrt{p}\zeta^{\frac{D-1}{2}})$ patterns in one iteration under a $\epsilon$-perturbation if
@@ 4.6 Extension of previous results for extra normalization step
We consider the image set induced by this transformation $\mathrm{im}(\hat{\bm{y}}_{\Psi}):=\left\{\hat{\bm{y}}_{\Psi}(\bm{z})\,:\,\bm{z}\in\mathbb{R}^{D}\right\}$. For $\ell_{2}$-normalization, $\hat{\bm{y}}_{\Psi}(\bm{z})=\frac{r\bm{z}}{\|\bm{z}\|}$, this image set is a $(D-1)\textsuperscript{th}$ sphere of radius $r$, and for layer normalization, $\hat{\bm{y}}_{\Psi}(\bm{z})=\mathrm{LayerNorm}(\bm{z};\eta,\bm{\delta})$ (14), it is the $(D-2)\textsuperscript{th}$-dimensional set $\left\{\bm{q}\,\,:\,\,\|\bm{q}-\bm{\delta}\|=\eta\sqrt{D}\,\,\wedge\,\,\bm{1}^{\top}(\bm{q}-\bm{\delta})=0\right\}$.
@@ 5.4 Guarantees for retrieval of pattern associations
For the sequential $k$-subsets problem in Example 2, we have r=2N1r=2N-1. Noting that any two distinct $\bm{y}$ and $\bm{y}^{\prime}$ differ in at most $2k$ variable nodes, and since each variable node can affect 6 bits (2 for $\bm{y}_{V}$ and 4 for $\bm{y}_{F}$), the Hamming distance between $\bm{y}$ and $\bm{y}^{\prime}$ is at most $12k$, therefore we have $D_{i}=\sqrt{12k}$, which leads to the condition $\Delta_{i}\geq\frac{6k}{\beta}+12Mk\epsilon$.
@@ 6.1 Free recall
Algorithm 1 Free recall with constrained sparsemax. $\bm{X}\in\mathbb{R}^{N\times D}$ represents the memory and $\bm{q}\in\mathbb{R}^{D}$ denotes the query, initialized as an arbitrary cue. $T$ is the number of inner associative Hopfield iterations. NN is the number of memory patterns, which equals the number of outer iterations during the recall process.
[…]
@@
5:    $\bm{\theta}\leftarrow\bm{X}\bm{q}$ $\triangleright$ Scores
@@
7:    $\bm{q}\leftarrow\bm{X}^{\top}\bm{p}$ $\triangleright$ Hopfield update
8:  $\bm{u}\leftarrow\bm{u}-\bm{p}$ $\triangleright$ Upper bound probabilities
@@
[…]
@@
4:  $\bm{p}\leftarrow\alpha\text{-entmax}\left(\beta\left(\bm{X}\bm{q}-\lambda\bm{a}\right)\right)$ $\triangleright$ Penalized probabilities
5:  $\bm{a}\leftarrow\tau\bm{p}+(1-\tau)\bm{a}$ $\triangleright$ Exponentially weighted average
6:  $\bm{q}\leftarrow\bm{X}^{\top}\bm{p}$ $\triangleright$ Outer Hopfield update
@@
9:    $\bm{q}\leftarrow\bm{X}^{\top}\bm{p}$ $\triangleright$ Inner Hopfield update